40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11207     Entry time: Wed Apr 8 03:44:48 2015
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Mediocre locking night 

It was only a mediocre locking night.  I was foiled for a long time by 2 things, both of which are now taken care of in the scripts, so I don't waste so much time again.

  • Somehow FM4 (LP700) in the CARM_A filter bank was turned off.  It took me a long time to figure this one out. 
    • At first, I had a 2056Hz oscillation, and then if I notched that, I would have problems at the edges of my notch.
    • I turned on the LP1000 in CARM_B (which we don't usually use) to fight the 2k resonances, but then I had violin mode problems. 
    • I couldn't turn off the violin mode filters on MC2 for the transition, so the edges of these notches were causing instability.
    • Anyhow, in the end, I realized that FM4 of CARM_B wasn't on, but it usually is.
    • It is now turned on in the carm_cm_up.sh script.
  • After that fiasco, I had trouble turning on the ASC loops.
    • Turns out we had left the offsets in place from last night in the ASC loops, so that when I zeroed the outputs of the transmission QPDs, the offsets in the ASC loops didn't make any sense, and they pushed the IFO severely out of alignment.
    • Now the ASC down script (which is run by the carm down script) zeros the filter bank offset values

Scripts are checked into the svn.


I used Q's handy-dandy 2D histogram plotter (..../scripts/general/dataHist, which I have taken the liberty of adding to the svn) to set the PRCL offset when I was locked on REFL165.  Here is a version of the plot, when I had an offset of +10 in the PRCL filter bank. There was so much noise on the PRCL input that I quit bothering to try and put in an excitation or ramp the offset value.  Note that I have since moved this offset to PRCL_A's offset instead, so taking this plot again should have PRCL_IN1 centered around zero.

I had trouble doing something similar for PRCL when I was locked on REFL55.  At first, the offset was so poor that POP110 was only about half the value it was when locked on REFL165, and it had a huge amount of RIN.  I tried just doing a z avg of the PRCL_B_IN1 (REFL55I) while locked on REFL165, and that said that REFL55I had an offset of +33.8 counts, so I tried an offset of -33.8 counts to get to zero.  But, that was still terrible for POP110 power.  As I increased the offset, eventually up to +30 counts, POP110 kept getting better and better.  I lost lock at that point (while trying to get 10 sec of histogram data), so I'm not sure that +30 is the final value.  I want to also get equivalent histograms with POP22 and POPDC (and maybe arm transmissions?) as the X-axis on these plots.  There's no excitation, so all of these can be collected at once.


By babysitting the ITM alignment (looking at the rough DC values of the optical lever error siganls), and doing a little adjustment of the ASC differential offset, I was able to keep lock a few times for more than 2 or 3 minutes while all 1f.  Not a whole lot longer than that though, even if I wasn't "poking" the interferometer other than maintaining alignment.

Attachment 1: PRCL_R165_offsetplus10.png  106 kB  | Hide | Hide all
PRCL_R165_offsetplus10.png
ELOG V3.1.3-